OpenStudy (anonymous):

write the function f(x) = cosh(2x)+2sinh^2x in exponential form

OpenStudy (anonymous):

\[ \cosh ( 2 x) = \frac { e^{2x} + e^{-2x}}2\\ \sinh ( 2 x) = \frac { e^{2x} - e^{-2x}}2\\ \]

OpenStudy (anonymous):

Add them up. It should be easy to find the answer.

OpenStudy (anonymous):

Its sinh^2(x) its square not 2x

OpenStudy (anonymous):

apply the same concept as shown then.

OpenStudy (anonymous):

\[ \cosh ^2( x) = \left ( \frac { e^{x} + e^{-x}}2 \right)^2\\ \sinh ^2 x = \left ( \frac { e^{x} - e^{-x}}2\right)^2\\ \]

OpenStudy (anonymous):

Does this mean that cos and sin functions are actually two functions combined?

OpenStudy (anonymous):

\[ \\sinh^2(x) =\frac {e^{4x} -2 + e^{-4x}}{4} \]

OpenStudy (anonymous):

\[ 2 \sinh^2(x) =\frac {e^{4x} -2 + e^{-4x}}{2} \]

OpenStudy (anonymous):

Now add cosh(2x) to the above and you will be done.

OpenStudy (anonymous):

Recall that \[\cosh ( 2 x) = \frac { e^{2x} + e^{-2x}}2\\ \]

OpenStudy (anonymous):

Add \[ \cosh ( 2 x) = \frac { e^{2x} + e^{-2x}}2\\ \text { to }\\ 2 \sinh^2(x) =\frac {e^{4x} -2 + e^{-4x}}{2} \] and you are done.